# Reducing Size Bias in Sampling Infectious Disease Spread on Networks

### v2

@@ 3.2.2 Average secondary infections
Table 3: U-test statistics (Hypotheses (2.5) and (2.6)) for comparing the estimates of the number of secondary infections in the samples generated using the RW and the MHRW sampling algorithms with the UN, with and without duplicate nodes.
@@
Results for the time-to-infection disease metric are described in C.1. We observed that in Figure 8, estimates from both sampling algorithms (with or without duplicates) are close to the UN. Although the RW samples slightly underestimate the time-to-infection, the MHRW samples align well with the UN.
@@
The RW samples overestimated the disease metrics regardless of whether duplicate nodes were removed or retained. The size bias in degree distribution (Section 3.1), for the RW samples, leads to overestimating the number of secondary infections for all networks, as the nodes with high degrees are more likely to become infected.
The MHRW samples align well with the UN for the ER and the SW networks, for all disease metrics, only when duplicate nodes are retained in the sample. The removal of duplicate nodes leads to an overestimation of the disease metrics (Section 3.2.1 and 3.2.2). One reason for overestimation after removing duplicates for MHRW samples is the reduction in the number of unique nodes in the MHRW samples, with a comparatively smaller reduction in the number of unique infected nodes. This may occur because the MHRW algorithm tends to get ‘stuck’ at nodes with very low connectivity due to fewer potential candidate nodes for the acceptance-rejection mechanism. Another contributing factor may be the relatively higher presence of highly connected nodes—which are more likely to become infected—in the MHRW samples compared to the UN, as observed in Figure 3.
Notably, for the SF network, both the sampling algorithms fail to obtain representative samples of the UN, resulting in high variation in estimates of the degree distribution and disease metrics. The SF network structure is complex compared to the ER and the SW networks; there is more heterogeneity in degree values. Also, the structure consists of two types of components: 1) core nodes of very high degree, and 2) peripheral nodes of low degree, which constitute the majority of the network. This diversity leads to dependency on the starting node for both sampling algorithms, leading to high variability in sample estimates.
@@
Furthermore, the performance of any sampling algorithm depends heavily on the complexity of the UN structure and disease parameters (such as transmission and recovery rates). For complex network structures, such as scale-free networks, sample estimates tend to be highly variable and sensitive to disease parameters, underscoring the importance of providing uncertainty intervals with estimates to policymakers and considering alternative sampling algorithms (Leskovec and Faloutsos, 2006a), depending on the availability of computational resources.
@@ 4 Cattle Movement Network
(a) A part of the cattle movement network.
(b) Degree distribution of cattle data.
Figure 6: Cattle movement network data.
In Section 3.2, we compared the network properties and disease metrics estimated from the RW and the MHRW sampling algorithms, on three theoretical network types. Now, we implement both sampling algorithms on movement network data. We use the cattle movement data provided by the British Cattle Movement Service (BCMS), which contains information about the birth, death, and movement of the cattle.
We use cattle movement data from January 2018 to March 2018, which is recorded at day level. We consider cattle holding places (farms, marketplaces, and slaughterhouses) as nodes and an edge is created between two nodes if there is cattle movement between them. Note that movement does not occur through all the edges every day, which implies that network structure changes with time. However, for this study, we assume that all edges are active and the network is a static network.
@@ 4.1 Data Description
The data contains movement details of $1,135,502$ cattle; there are two movement records on average, per cattle. There are $46,512$ holding places, which are the nodes of the network. A static and undirected network is created with $46,512$ nodes and $159,036$ edges. Note that an edge between a pair of nodes is counted only once, in case of multiple movement records. A part of the movement network is shown in Figure 6(a); where we see a mix of large connected components, small clusters of nodes and isolated node pairs.
@@
Figure 7: The ratio of disease metric estimates between sampled generated using RW and MHRW algorithm, relative to the UN for two scenarios: (1) samples with duplicate nodes, and (2) samples without duplicate nodes. The light red and blue bands indicate the range of one standard deviation. disease metric estimates are based on $500$ SIR model simulations on cattle movement network with $\gamma=1$. Sample estimates are derived from $10,000$ samples (size $2500$) for cattle movement network using RW and MHRW.
The network’s degree ranges from $1$ to $6,556$, with an average degree of $6$. Notably, $30\%$ of the nodes have a degree of $1$, while only $2.69\%$ of the nodes have a degree greater than $20$. There is only one node with $6,556$ edges, which would be a marketplace.
The degree distribution for the network is shown in Figure 6(b), up to a degree value of $20$. The degree distribution of the cattle network is fitted with an exponential distribution (Balakrishnan, 2019) with rate parameter $0.4$ and location parameter $1$ (the distribution is shifted by one unit to the right along the horizontal axis), represented by the solid black line in Figure 6(b). Other studies state that cattle movement network follows a power-law degree distribution (Christley et al., 2005, Fielding et al., 2019). One reason for this difference is the amount and period of data, we have only three months of data (from 2018) and in these studies, data for multiple years before 2018 is analysed (Duncan et al., 2022).
We observed the presence of clusters with few connections outside the cluster, which is similar to the structure of the SW networks and agrees with the results in (Fielding et al., 2019). We also observe a uniform connectivity pattern (average degree $2$) similar to the SW and ER networks if we exclude the nodes with one connection or with degree $>1000$. We observe the SF network’s trait of the presence of “hubs” in the cattle movement network, which is suggested by the presence of $0.06\%$ nodes with degree $\geq 1000$ and $30\%$ nodes with degree $1$. There are a few nodes with very high degrees, likely markets, facilitating extensive animal movement and which can potentially act as key nodes for disease spread.
@@ 4.2 Results
Due to a lack of disease data, we study a hypothetical disease spread scenario, for example, bovine tuberculosis (Bekara et al., 2014), we run $500$ simulations of the SIR model on the cattle movement network using the method and parameters from Section 2.4. The RW and the MHRW sampling algorithms are used to generate $10,000$ samples with $2500$ as sample size ($5\%$ of the network size). The number of samples, sample size, and number of SIR model simulations are chosen to adhere precision to the 2nd decimal place as in Section 3.
@@
In Figure 7, we compare three disease metrics in RW and MHRW samples relative to the cattle network, as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for the cattle movement network. The three disease metrics are: 1) proportion of infected nodes, 2) number of secondary infections, and 3) time-to-infection. Panel 1 (Figure 6(c), 6(d), 6(e)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 6(f), 6(g), 6(h)) shows those without duplicates. The closer the disease metric estimates made from the samples are to the solid horizontal black line, the closer they are to the underlying network.
@@
We observe that the estimate of the proportion of infected nodes in an epidemic is overestimated by RW samples in both panels (with and without duplicate nodes); see Figure 6(c) and Figure 6(f). However, the estimate from MHRW samples aligns very well with cattle network when duplicate nodes are retained (Figure 6(c)), but it overestimates when duplicate nodes are removed (Figure 6(f)). In the case of samples without duplicate nodes, the estimates of the RW and the MHRW samples are the same for $\beta>0.3$.
@@
The RW samples greatly overestimate the number of secondary infections when duplicate nodes are retained compared to the MHRW samples, which align well with the cattle network (Figure 6(d)). We observe a significant difference in the estimates when duplicate nodes are removed. The RW sample still overestimates but it is in the range of $10-20$, whereas when duplicates are retained the range is $40-130$. However, without duplicate nodes, the MHRW samples overestimate the number of secondary infections by approximately $7$ times. Also, the variability in the estimates from the MHRW samples is higher when duplicate nodes are removed compared to RW samples.
@@
In Panel 1 of Figure 7, the estimate of time-to-infection is overestimated by the MHRW samples and underestimated by the RW samples. The variability in estimates is extremely high for the MHRW samples compared to the thin band for the RW sample estimates (Figure 6(e)). Interestingly, after removing the duplicate nodes, both sampling algorithms underestimate the time to get the infection, which deviates further from the cattle network estimate as the $\beta$ value increases.
@@
Figure 7 indicates that the estimates from MHRW samples are closer to the cattle network than the RW samples. However, the variability is significantly higher for the MHRW samples for the number of secondary infections and the time-to-infection.
@@
The trend of disease metric estimates on the cattle movement network is consistent with those seen in the ER and SW networks. The RW samples provide conservative estimates of disease spread, whereas the MHRW samples yield estimates closer to the underlying network when duplicate samples are retained. For highly contagious diseases like Foot and Mouth Disease, which has a mortality rate of $20\%$ in young calves (World Organisation for Animal Health, ), the RW sampling algorithm is recommended. Conversely, for diseases such as Bovine Viral Diarrhea, with a mortality rate of approximately $7\%$ (Dobos et al., 2024), the MHRW algorithm is preferable as it provides estimates more representative of the underlying network, which is particularly valuable for optimizing intervention policies.
5 Conclusions
@@
In this study, we assessed the performance of the Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW) sampling algorithms for estimating disease metrics across three network types: Erdős-Rényi (ER), Small-World (SW), and Scale-Free (SF). These network types were chosen because they capture essential properties of real-world networks. For instance, sexual contact networks often resemble ER networks (Holme, 2013), while the SARS-CoV-2 transmission network in Houston, Texas, has been shown to align more closely with SF networks (Fujimoto et al., 2023). Using a stochastic SIR model, we examined how size bias, caused by the over-representation of highly connected individuals, in samples generated using the RW algorithm, leads to biased estimates of disease spread. Additionally, we investigated how the inclusion of duplicate nodes influences the estimation of key disease metrics. Our analysis focused on three key disease metrics: the proportion of infected individuals, the average number of secondary infections, and the time-to-infection.
@@
Our findings (Section 3.2) highlight the suitability of the MHRW algorithm for ER and SW networks. By reducing size bias, MHRW samples consistently produced estimates that closely aligned with the underlying network’s estimates, particularly when duplicate sampled nodes were retained. In contrast, the RW algorithm produced estimates with significant overestimation, approximately $25\%$, for the proportion of infected individuals and secondary infections. Hence, the RW algorithm is less reliable for precise estimates but can be valuable for early interventions in high-fatality diseases like COVID-19 and Ebola. Conversely, the MHRW algorithm is more appropriate for diseases, like seasonal influenza. Notably, estimates of the time-to-infection metric align closely with the underlying network for both sampling algorithms.
@@
In SF networks, the inherent heterogeneity and presence of high-degree hubs introduced substantial variability in the estimates derived from both the RW and MHRW samples. The high variability in the estimated disease metrics suggests that neither method is optimal for SF networks. Instead, alternative approaches, such as separate sampling of core and peripheral nodes using algorithms like the Sampling for Large-Scale networks at low Sampling Rates (SLSR) (Jiao, 2024), may be required for more representative sampling.
@@
In Section 4, we analyzed a cattle movement network that exhibits characteristics of ER, SW, and SF networks, this further supported the findings above. MHRW sample estimates were more representative of the underlying network when duplicate nodes were retained, while RW samples significantly overestimated the proportion of the infected farms by approximately $100\%$ and the number of secondary infections by over $900\%$, reflecting size bias. Also, RW samples underestimated the time-to-infection by approximately $40\%$. These findings suggest that RW’s conservative estimates are advantageous for rapid containment measures in highly contagious diseases, such as Foot-and-Mouth Disease. Conversely, the more precise estimates from MHRW sampling are better suited for resource-efficient intervention strategies in less severe diseases, such as Bovine Viral Diarrhea.
@@
Our analysis revealed that the sensitivity of both sampling algorithms to recovery rates significantly impacts the accuracy of disease metric estimates (C.2). While the difference in the proportion of infected nodes estimated by the two sampling algorithms reduces with the recovery rate, variability in the estimates of secondary infections and time-to-infection increases with higher transmission rates.
@@
In future works, to further understand the subtle differences between the RW and MHRW sampling algorithms with respect to network structure, it would be valuable to explore the impact of selecting a specific starting node, rather than a random node, particularly for SF networks. As observed in Figure 9, the difference in disease estimates between the RW and MHRW samples decreases with the recovery rate. It would be interesting to determine the value of the recovery rate at which this change starts, as it could guide the choice of sampling algorithm. Additionally, exploring the sensitivity of these estimates to variations in network parameters, such as network size, average degree, and the power-law exponent for SF networks, would provide further insight.

### v3

@@ 3.2.2 Average secondary infections
Table 3: U-test statistics (Hypotheses 2.6 and 2.7) for comparing the estimates of the number of secondary infections in the samples generated using the RW and the MHRW sampling algorithms with the UN, with and without duplicate nodes.
3.3 Time-to-infection
Panel 1: Samples with duplicate nodes.
(a) Erdős-Rényi network
(b) Small-world network
(c) Scale-free network
Panel 2: Samples without duplicate nodes.
(d) Erdős-Rényi network
(e) Small-world network
(f) Scale-free network
Figure 6: The ratio of the time-to-infection during an epidemic between RW (dashed red) and MHRW (dotted blue) samples, relative to their respective UNs, is shown for (1) retained and (2) removed duplicate nodes. The light red and blue bands indicate the range of one standard deviation. The UN estimates are based on SIR model simulations ($\gamma=1$) for $10,000$ networks of each type (the parameters are found in Table 1) generated with the NetworkX library (Hagberg et al., 2008). Sample estimates are derived from $100$ samples (size $500$) for $10,000$ networks using RW and MHRW.
Network Sample With duplicate nodes Without duplicate nodes
RW MHRW RW MHRW
ER $6.2\times 10^{11}$ $5.98\times 10^{11}$ $6.2\times 10^{11}$ $6.08\times 10^{11}$
SW $6.1\times 10^{11}$ $6.0\times 10^{11}$ $3.91\times 10^{11}$ $4.51\times 10^{11}$
SF $1.26\times 10^{11}$ $7.37\times 10^{10}$ $8.45\times 10^{10}$ $5.99\times 10^{10}$
Table 4: U-test statistics (Hypotheses 2.8 and 2.9) for comparing the estimates of the time-to-infection of the UN with the samples generated using the RW and the MHRW sampling algorithms, with and without duplicate nodes.
In Figure 6, we compare estimates of the time-to-infection from the RW and MHRW samples with the UN for increasing $\beta$ values and $\gamma=1$ for the three networks (ER, SW, and SF). In Panel 1 (Figure 5(a), 5(b), 5(c)) results for samples with duplicate nodes and in Panel 2 (Figure 5(d), 5(e), 5(f)) results for samples without duplicate nodes are shown.
We observe that for ER networks and $\beta<0.2$, there is a large variation in the estimates of both the sampling algorithms in both panels. For $\beta\geq 0.2$, the MHRW sample estimates align well with the UN, while the RW samples slightly underestimate the time-to-infection if duplicate nodes are retained in the samples (Figure 5(a)). When duplicate nodes are removed, the MHRW sample estimates slightly underestimate the time-to-infection (Figure 5(d)), although the behaviour of the RW sample estimates remains consistent.
For SW networks, a behaviour similar to that of the ER networks is observed (Figure 5(b), 5(d)). However, the large variation in the estimates for both sampling algorithms persists up to $\beta=0.5$ for the SW networks.
Interestingly, for SF networks, as shown in Figures 5(c) and 5(f), the behaviour of both algorithms remains consistent across both panels, i.e., estimates from both the algorithms align well with the UN and have a thin variation band. Overall, both sampling algorithms perform well across network types, particularly for the Scale-free (SF) networks, which is not the case for the other two disease metrics (the proportion of infected nodes and the number of secondary infections).
We conduct hypothesis testing to statistically compare the estimates from the sampling algorithms. First, we conduct the one-sample KS test to assess whether the time-to-infection follows a Normal distribution, as stated in Hypothesis 2.3, with significance level $\alpha=0.05$, for the ER, SW, and SF networks individually. The KS test statistic is summarized in B Table 7 for three data groups that are UN, RW, and MHRW for ER, SW, and SF networks. The KS statistic ranges from $0.5$ to approximately $0.7$ across networks and data groups, with p-values less than $0.05$ individually for all three networks. These results indicate significant deviation from normality, providing evidence to reject $H_{0}$ in Hypothesis 2.3.
Since time-to-infection does not follow a Normal distribution, we conduct a one-tailed U-test to compare the RW and the MHRW sample estimates with the UN individually for the ER, SW, and SF networks. We test two hypotheses, one for statistically significant overestimation (Hypothesis 2.8) and another for statistically significant underestimation (Hypothesis 2.9), with significance level $\alpha=0.05$. The statistics of the U-test are summarized in Table 4 for samples with duplicate nodes and samples without duplicate nodes.
The p-values are less than $0.05$ individually for three network types, for the RW samples (with and without duplicates), and for Hypothesis 2.9. This provides evidence to reject $H_{0}$ in Hypothesis 2.9 and indicates a significant underestimation of the time-to-infection.
For the MHRW samples, without duplicate nodes individually for all three networks and with duplicate nodes for the SW and SF networks, p-values are less than $0.05$, evidence to reject $H_{0}$ in Hypothesis 2.9, implying a significant underestimation of the time-to-infection. However, for the ER network, the p-value is less than $0.05$ (accept $H_{0}$ in Hypothesis 2.8) for the MHRW sample estimates with duplicate nodes, indicating that the MHRW significantly overestimates the time-to-infection.
In the case of time-to-infection, the U-test statistics (see Table 4) are similar for the RW sample estimates (with or without duplicate nodes) compared to the MHRW for the ER and SW networks, implying that both RW and MHRW estimates are closer to the UN. However, for SF networks, the U-test statistics are higher for RW sample estimates (with or without duplicate nodes) compared to the MHRW, implying that RW estimates are closer to the UN.
@@ 3.3 Time-to-infection
We observed that in Figure 6, estimates from both sampling algorithms (with or without duplicates) are close to the UN. However, on average, the RW samples underestimate and the MHRW samples overestimate the time-to-infection across three network types.
@@
The RW samples overestimated the proportion of infected nodes and the number of secondary infections and underestimated the time-to-infection regardless of whether duplicate nodes were removed or retained. The size bias in degree distribution (Section 3.1) for the RW samples leads to overestimating the number of secondary infections for all networks, as the nodes with high degrees are more likely to become infected.
The MHRW samples align well with the UN for the ER and the SW networks for all disease metrics only when duplicate nodes are retained in the sample. The removal of duplicate nodes leads to an overestimation of the disease metrics (Section 3.2.1, 3.2.2 and 3.3). One reason for overestimation after removing duplicates for MHRW samples is the reduction in the number of unique nodes in the MHRW samples, with a comparatively smaller reduction in the number of unique infected nodes. This may occur because the MHRW algorithm tends to get ‘stuck’ at nodes with very low connectivity due to fewer potential candidate nodes for the acceptance-rejection mechanism. Another contributing factor may be the relatively higher presence of highly connected nodes—which are more likely to become infected—in the MHRW samples compared to the UN, as observed in Figure 3.
Notably, for the SF network, both the sampling algorithms fail to obtain representative samples of the UN, resulting in high variation in estimates of the degree distribution and disease metrics. The SF network structure is complex compared to the ER and the SW networks; there is more heterogeneity in degree values. Also, the structure consists of two types of components: 1) core nodes of very high degree and 2) peripheral nodes of low degree, which constitute the majority of the network. This diversity leads to dependency on the starting node for both sampling algorithms, leading to high variability in sample estimates.
@@
Furthermore, the performance of any sampling algorithm depends heavily on the complexity of the UN structure and disease parameters (such as transmission and recovery rates). For complex network structures, such as SF networks, sample estimates tend to be highly variable and sensitive to disease parameters, indicating the importance of providing uncertainty intervals with estimates to policymakers and considering alternative sampling algorithms (Leskovec and Faloutsos, 2006a), depending on the availability of computational resources.
@@ 4 Cattle Movement Network
(a) Subset of the cattle movement network.
(b) Degree distribution.
Figure 7: Illustration of the cattle movement network data from BCMS.(a) A subset of the cattle movement network, where nodes represent individual farms and edges indicate recorded cattle movements between them.(b) Degree distribution of the cattle network. The red histogram represents the empirical data and the black curve shows an exponential fit for comparison.
In Section 3.2, we compared the network properties and disease metrics estimated from the RW and the MHRW sampling algorithms on three theoretical network types. Now, we implement both sampling algorithms on cattle movement network data. We utilise the cattle movement data provided by the British Cattle Movement Service (BCMS), which contains information on the birth, death, and movement of cattle.
We use cattle movement data from January 2018 to March 2018, recorded at day level. We consider cattle holding places (farms, marketplaces, and slaughterhouses) as nodes and an edge is created between two nodes if there is cattle movement between them. Note that movement does not occur through all edges every day, which implies that the network structure changes over time. However, for this study, we assume that all edges are active and that the network is static.
@@ 4.1 Data description
The data contains movement details of $1,135,502$ cattle; there are two movement records on average per cattle. There are $46,512$ holding places, which are the nodes of the network. A static and undirected network is created with $46,512$ nodes and $159,036$ edges. Note that an edge between a pair of nodes is counted only once in case of multiple movement records. A part of the movement network is shown in Figure 7(a); we observe a mix of large, connected components, small clusters of nodes, and isolated node pairs.
@@
Figure 8: The ratio of disease metric estimates between sampled generated using RW and MHRW algorithm, relative to the UN for two scenarios: (1) samples with duplicate nodes, and (2) samples without duplicate nodes. The light red and blue bands indicate the range of one standard deviation. disease metric estimates are based on $500$ SIR model simulations on the cattle movement network with $\gamma=1$. Sample estimates are derived from $10,000$ samples (size $2500$) for cattle movement network, using RW and MHRW.
The network’s degree ranges from $1$ to $6,556$, with an average degree of $6$. Notably, $30\%$ of the nodes have a degree of $1$, while only $2.69\%$ of the nodes have a degree greater than $20$. There is only one node with $6,556$ edges, which should be a marketplace.
The degree distribution for the network is shown in Figure 7(b), up to a degree value of $20$. The degree distribution of the cattle network is fitted with an exponential distribution (Balakrishnan, 2019) with rate parameter $0.4$ and location parameter $1$ (the distribution is shifted by one unit to the right along the horizontal axis), represented by the solid black line in Figure 7(b). Other studies state that cattle movement network follows a power-law degree distribution (Christley et al., 2005, Fielding et al., 2019). One reason for this difference is the amount and period of data; we have only three months of data (from 2018), whereas these studies analyse data for multiple years preceding 2018 (Duncan et al., 2022).
We observed the presence of clusters with few connections outside the cluster, which is similar to the structure of the SW networks and agrees with the results in (Fielding et al., 2019). We also observe a uniform connectivity pattern (average degree $2$), similar to the SW and ER networks, if we exclude nodes with one connection or a degree greater than $1000$. We observe the SF network’s trait of the presence of “hubs” in the cattle movement network, which is suggested by the presence of $0.06\%$ nodes with degree $\geq 1000$ and $30\%$ nodes with degree $1$. There are a few nodes with very high degrees, likely markets, that facilitate extensive animal movement and can potentially act as key nodes for disease spread.
@@ 4.2 Results
Due to a lack of disease data, we study a hypothetical disease spread scenario, for example, bovine tuberculosis (Bekara et al., 2014). We run $500$ simulations of the stochastic SIR model on the cattle movement network using the methodology and parameters from Section 2.4. The RW and MHRW sampling algorithms are used to generate $10,000$ samples, with a sample size of $2,500$ ($5\%$ of the network size). The number of samples, sample size, and number of SIR model simulations are chosen to be accurate to the second decimal place, as in Section 3.
@@
In Figure 8, we compare three disease metrics in RW and MHRW samples relative to the cattle network, as the transmission rate $\beta$ increases, with a recovery rate $\gamma=1$, for the cattle movement network. The three disease metrics are: 1) proportion of infected nodes, 2) number of secondary infections, and 3) time-to-infection. Panel 1 (Figure 7(c), 7(d), 7(e)) shows the results for samples with duplicate nodes, while Panel 2 (Figure 7(f), 7(g), 7(h)) shows those without duplicates. The closer the disease metric estimates are to the solid horizontal black line, the better the agreement to the underlying network metric values.
@@
We observe that the estimate of the proportion of infected nodes in an epidemic is overestimated by RW samples in both panels (with and without duplicate nodes); see Figure 7(c) and Figure 7(f). However, the estimate from MHRW samples aligns very well with the cattle network when duplicate nodes are retained (Figure 7(c)), but it overestimates when duplicate nodes are removed (Figure 7(f)). In the case of samples without duplicate nodes, the estimates of the RW and the MHRW samples are the same for $\beta>0.3$.
@@
The RW samples overestimate the number of secondary infections when duplicate nodes are retained compared to the MHRW samples for the ER and SW networks, which align well with the cattle network (Figure 7(d)). We observe a significant difference in the estimates when duplicate nodes are removed. The RW sample overestimates the number of secondary infections by a factor of 10–20, which increases to a factor of 40–130 when duplicate nodes are retained. In comparison, the MHRW sample without duplicates overestimates by a factor of approximately 7. Also, the variability in the estimates from the MHRW samples is higher when duplicate nodes are removed compared to RW samples.
@@
In Panel 1 of Figure 8, the estimate of time-to-infection is overestimated by the MHRW samples and underestimated by the RW samples. The variability in estimates is extremely high for the MHRW samples compared to the thin band for the RW sample estimates (Figure 7(e)). Interestingly, after removing the duplicate nodes, both sampling algorithms underestimate the time-to-infection. This estimate deviates further below the cattle network estimate as the $\beta$ value increases.
@@
Figure 8 indicates that the estimates from MHRW samples are closer to the cattle network than the RW samples. However, the variability is significantly higher for the MHRW samples for the number of secondary infections and the time-to-infection compared to the RW samples.
@@
The trend of disease metric estimates on the cattle movement network is consistent with those seen in the ER and SW networks. The RW samples provide conservative estimates of disease spread, whereas the MHRW samples yield estimates closer to the UN when duplicate samples are retained. For highly contagious diseases, such as Foot and Mouth Disease, which has a mortality rate of $20\%$ in young calves (World Organisation for Animal Health, 2024) and an average of $20$ secondary infections for direct transmission (Paton et al., 2018), the RW sampling algorithm is recommended. Conversely, for diseases such as Bovine Viral Diarrhoea, with a mortality rate of approximately $7\%$ (Dobos et al., 2024) and the average number of secondary infections as $0.36$ (Isoda et al., 2025), the MHRW algorithm is preferable as it provides more representative estimates, which is particularly valuable for optimizing intervention policies.
5 Conclusions and Further Directions
@@
In this study, we assessed the performance of the Random Walk (RW) and Metropolis-Hastings Random Walk (MHRW) sampling algorithms in estimating disease transmission metrics across three network types: Erdős-Rényi (ER), Small-world (SW), and Scale-Free (SF). We chose these network types because they capture essential properties of real-world networks. For instance, sexual contact networks often resemble ER networks (Holme, 2013), the SARS-CoV-2 transmission network in Houston, Texas, has been shown to align more closely with SF networks (Fujimoto et al., 2023), and the rabies transmission network of Serengeti Lions is observed to be similar to the SW networks (Craft et al., 2011). Using a stochastic Susceptible-Infected-Recovered (SIR) model and the RW algorithm, we examined how size bias, the over-representation of highly connected individuals in samples, leads to biased estimates of disease spread. Our analysis focused on three key disease metrics: the proportion of infected individuals (epidemic size), the average number of secondary infections (effective reproduction number), and the time-to-infection. Additionally, we investigated how the inclusion of duplicate nodes influences the estimation of key disease metrics.
@@
Our findings in Section 3.2 demonstrate that the MHRW algorithm is suitable for ER and SW networks. The RW algorithm produced estimates with significant overestimation, approximately $25\%$, for the proportion of infected individuals and secondary infections. In contrast, by reducing size bias, MHRW samples consistently produced estimates of the three disease metrics that closely aligned with the underlying network’s estimates, particularly when duplicate sampled nodes were retained. Hence, the RW algorithm is less reliable for precise estimates but can be valuable for early interventions in high-fatality and/or fast-spreading epidemics, such as COVID-19 and Ebola. Conversely, the MHRW algorithm is more suitable for slower and lower-severity epidemics, such as seasonal influenza. Notably, for both sampling algorithms, estimates of time-to-infection align closely with the underlying network for all three network types.
@@
In SF networks, the inherent heterogeneity and presence of high-degree hubs introduced substantial variability in the estimates derived from both the RW and MHRW samples. The high variability in the estimated disease metrics suggests that neither method is optimal for SF networks. Alternative approaches, such as separate sampling of core and peripheral nodes using algorithms like the Sampling for Large-Scale Networks at Low Sampling Rates (SLSR) (Jiao, 2024), may be required for more representative sampling.
@@
In Section 4, we analysed cattle movement network data that exhibit hybrid characteristics from ER, SW, and SF networks; this further supported the findings above. MHRW sample estimates were more representative of the underlying network when duplicate nodes were retained, while RW samples significantly overestimated the proportion of the infected farms by approximately $100\%$ and the number of secondary infections by over $900\%$ due to size bias. Also, RW samples underestimated the time-to-infection by approximately $40\%$. These findings suggest that RW’s conservative estimates are suitable for rapid interventions in highly contagious diseases, such as the Foot-and-Mouth disease. Conversely, the more precise estimates from MHRW sampling are better suited for resource-efficient intervention strategies in less severe and slower epidemics, such as Bovine Viral Diarrhoea.
@@
Our analysis revealed that the sensitivity of both sampling algorithms to recovery rates significantly impacts the accuracy of disease metric estimates (C.1). The difference in the proportion of infected nodes estimated by the two sampling algorithms decreases with an increase in the recovery rate.
In this study, we have demonstrated the effectiveness of the MHRW algorithm over the RW algorithm for reducing size bias in estimating disease metrics for three key network types. However, it is essential to note that MHRW, despite its advantages in reducing size bias, has limitations in networks with high-degree variance, where slow mixing can lead to non-representative samples, such as SF networks. Additionally, MHRW relies on the assumption that node degrees are known or accessible during sampling, which may not always be feasible in dynamic networks. While retaining duplicate nodes improves the accuracy of the MHRW estimates, it increases computational costs and may reduce the scalability of the method in larger, more complex networks.
The use of the Gillespie algorithm for simulating stochastic SIR disease dynamics also introduces certain constraints; it can become computationally expensive for larger networks, limiting the ability to perform large-scale simulations. Additionally, the choice of parameter values for transmission and recovery rates may not fully capture the diversity of epidemic dynamics in different settings. Although we selected parameters that reflect realistic disease scenarios, the variability in disease dynamics across different populations and network structures may affect the generalizability of our results. To account for variability in network topology, we generated $10,000$ networks of each type, considering specific values for the number of nodes, average degree, edge creation probability, and power-law exponent. However, these network parameters do not cover the full spectrum of all possible topologies within each network family. Consequently, further research is necessary to assess how variations in network structure and epidemiological parameters influence the robustness of our findings.
@@
Despite these limitations, our findings provide valuable insights into the behaviour of the RW and MHRW sampling algorithms and suggest areas for further work. To further understand the subtle differences between the RW and MHRW sampling algorithms with respect to network structure, it would be valuable to explore the impact of selecting a specific starting node rather than a random node, particularly for SF networks. Furthermore, as observed in Figure 9, the difference in disease estimates between the RW and MHRW samples decreases with the recovery rate. It would be interesting to determine the value of the recovery rate at which this change begins, as it could guide the choice of sampling algorithm. Additionally, exploring the sensitivity of these estimates to variations in network parameters, such as network size, average degree, and for SF networks, the power-law exponent, would provide further insights.
To provide deeper insights into the mechanisms driving size bias in sampling, one could study a configuration model with a Negative Binomial (NB) degree distribution, which allows systematic variation of the dispersion parameter to interpolate between Poisson-like (light-tailed) and heavy-tailed (power-law) degree distributions. We explored this approach and found several important subtleties. Notably, when the dispersion parameter $(r)$ is large, for example, $r=100$, the NB degree distribution closely resembles that of ER networks, characterised by homogeneous connectivity and a few high-degree nodes. This overlap arises because both NB and binomial distributions converge toward a Poisson distribution under specific conditions. On the other hand, decreasing the dispersion parameter increases the variance of the NB degree distribution, yielding heavier tails and greater heterogeneity. However, maintaining an average degree of $5$, as chosen in this study, with low dispersion values (for example, $r=0.001$ or $0.01$) results in a large number of nodes becoming isolated, significantly reducing the size of the connected component and, consequently, comparability. For example, networks generated using $r=0.01$ with $10,000$ nodes and average degree $5$ result in fewer than $1,000$ connected nodes, rendering them unsuitable for comparison with the ER, SW, and SF networks used in this study. Thus, although NB-based configuration models offer a promising direction for understanding size bias in relation to network heterogeneity, careful consideration of dispersion effects on network connectivity and comparability is essential. We recommend future work to explore this space further, potentially relaxing constraints on the average degree or network size to investigate size bias under more extreme connectivity profiles.